# Network Intrusion Datasets: A Survey, Limitations, and Recommendations

### v2

@@ Abstract
[inst2] organization=Kempelen Institute of Intelligent Technologies, addressline=Bottova 2/A, city=Bratislava, postcode=811 09, country=Slovakia
@@ 1 Introduction
The only truly secure system is one that is powered off, cast in a block of concrete and sealed in a lead-lined room with armed guards – and even then I have my doubts.” [1]. The legendary quote by Eugene H. Spafford from 1989 is not any less relevant nowadays. Each year, cyberattacks like malware or denial of service (DoS) cause losses of trillions of USD [2]. With recent attacks on hospitals and medical devices, even people’s lives are in danger [3]. Therefore, robust security is required to minimize an attack surface and mitigate potential harm.
@@
An Intrusion Detection System (IDS) is a device or software application monitoring the activity of a target system or a computer network to identify their potential unauthorized use, misuse, or abuse [5]. Depending on the attack detection paradigm, we distinguish between signature-based (react on known attack patterns) and anomaly-based (react on unknown deviation from the norm) systems. Such systems can operate on the host level – Host-based IDS (HIDS) or the network – Network-based IDS (NIDS). Additional IDSs characteristics include the ability to operate in real-time and whether they merely report the attack or actively participate in its mitigation [6, 7].
@@
Despite the dire need for data, a general consensus within the network intrusion detection (NID) domain over the past decade was that there had been a lack of quality datasets for evaluating NIDS proposals [10, 11, 12]. This factor has significantly contributed to the poor adoption of ML-based methods into real-world detection systems despite huge scientific efforts [13].
@@
As we show in this work, it is questionable whether data scarcity is still the main issue. Instead, we argue that datasets exist, but the community is largely unaware of their existence, as well as the best practices for their creation and usage. This can be primarily attributed to non-standardized NID data repositories and neglected domain-specific properties that must be considered when handling the data. As a result, many studies present biased evaluations and over-optimistic results [15].
@@
    In order to aid in data selection, we extract 13 properties for each dataset by manually downloading and analyzing the data. We also study the popularity of datasets in state-of-the-art NIDS research. To enhance accessibility, we share our analysis notebooks on GitHub [^1] https://github.com/xGoldy/nid-datasets, enabling others to review the data structure without downloading it.
@@ 2 Survey Methodology
In order to ensure comprehensive, unbiased, and reproducible results, we followed the Systematic Literature Review (SLR) methodology for computer science [16] to conduct our survey. This section outlines the survey’s scope and research questions, while further details are provided in A.
@@ 3. RQ3:
    Regarding the data properties from RQ2, are there any prevalent patterns or trends in the dataset creation?
@@ 3 Domain-Specific Properties and Their Impact on Data
The research on intrusion detection (ID) initially revolved around static signatures or simple statistical methods. Nevertheless, these methods require significant manual interventions, such as maintaining signature databases and threshold values. Motivated to minimize the need for manual labor and scale to big data more efficiently, the research community has gradually adopted data-driven approaches for ID. Nowadays, AI, and especially ML, are crucial parts of most ID research [17].
@@ 3.1 When Hell Comes to Earth: Unpleasant Properties of the Network Intrusion Detection Domain
Network intrusion detection combines properties of both cyberthreat detection and computer networking domains. These properties significantly influence how the data are collected, handled, and interpreted. If machine learning techniques are used for detection, an assumption on the data to be independent and identically distributed (i. i. d.) is also introduced [19].
In essence, the i. i. d. assumes the data to be uncorrelated, while the data for training a model to be similar, i.e., drawn from the same distribution, as the future observed data. If not met, the model’s performance might be negatively affected. However, the intrinsic properties of cyberthreat detection and computer networking domains interfere with this assumption and thus impair ML deployment in real-world scenarios [17]. In the following paragraphs, we dive deeper into these properties and outline their potential impacts, as summarized in Figure 1.
@@
    Adversarial environment: In addition to natural networks’ dynamicity, the NID domain has to assume the implicit presence of adversaries. Due to the endless cybersecurity arms race, attacks continuously evolve, and new vulnerabilities are dicovered. Moreover, the attackers may also attempt to bypass detection via adversarial samples crafted with obfuscation techniques or tiny perturbations of the input data [21].
@@
    High cost of errors: Unlike other ML applications, the relative cost of any output error (i.e., misclassification) is very high. Regular applications can optimize for lower false alarm (false positive) or higher detection (true positive) rates. However, in NID, both need to be considered. Even a single missed anomaly might lead to a successful intrusion with dire consequences. On the other hand, a low relative false alarm rate can still create a tremendous amount of actual alarms due to a huge traffic imbalance, known as base-rate fallacy [22]. For instance, with a 99% true positive and 1% false positive rate in a 1:100 class balance, we receive 100 false alarms for every 99 true attacks.
@@
    Data confidentiality: Captures from real networks might contain highly sensitive and confidential information, such as IP addresses or packet payloads, significantly limiting their sharing. Although anonymization techniques [28] like IP address remapping or payload stripping can be used, they decrease the data’s value – e.g., payload analyzers NIDSs cannot be used.
@@ 3.2 Limitations of Existing Datasets
    Testbed limitations: Due to the limitations of real-world data, most public datasets contain emulated traffic captured in a lab testbed environment. Despite some benefits of such an approach, its data are not guaranteed to resemble realistic network behavior and threats [36]. Such datasets tend to model over-simplified network conditions, failing to capture their real-world complexity and diversity [33, 8]. The lack of realism in benchmark data is considered a major obstacle in deploying anomaly-based NIDSs in real-world production environments [13, 32]. For instance, Catillo et al. [37] have shown that many attacks from DoS public datasets do not resemble real-world characteristics due to being inefficient (not achieving their malicious intent) or assuming non-realistic environment configurations.
@@
    Reproducibility: In computer networking, it is infeasible to collect exactly the same data twice, even in identical environments. This property disallows the reproducibility of the raw packet data. Although it would not be a critical limitation by itself, reproducibility is further restricted for datasets publishing only extracted features (e.g., network flows) without packet data. Even if some datasets do provide raw data, they rarely publish source code and configuration for feature extraction and labeling. In all cases, reproducing the dataset (or even extracting the same features from packets) is infeasible, disallowing data validation and its possible extensions in future research [45].
@@
    Incorrect labeling: Due to networking dynamicity and the common methods for traffic labeling (based on the attack time and the source and destination addresses of the attacker/victim [47]), entries in NIDS datasets are prone to mislabeling. The issue has been demonstrated by Liu et al. [48] and Flood et al. [8] through analyses of several popular benchmark datasets. Such mislabeling can result in unrealistic results and benchmark destabilization [49].
@@
    Methodically wrong data handling: In addition to mislabeling, several other factors, primarily originating from human error in data collection and processing, affect the final dataset. A common issue we encountered relates to improper timestamping, as some datasets contain incorrect timestamps or time inconsistencies of events (e.g., attacks) with the documentation. We also observed disordered packets, undocumented capture gaps, and capture interruptions, resulting in improperly terminated files. Data processing issues, often due to bugs in feature extractor software [33], include missing values, duplicated features or entries, and zero-variance features (e.g., all zeros). These issues are frequently unreported, creating ambiguity for users and undermining the soundness of their conclusions.
@@ 4 Network Intrusion Data and Its Properties
Depending on the dataset goal and capture attributes, we characterize network intrusion data by various properties that influence their use cases and the soundness of methods utilizing them. In this survey, we collected 13 essential properties to provide an unbiased, comprehensive view and comparison of available datasets for NID purposes. A portion of them was adopted from Ring et al.’s survey [14].
@@ 4.1 General Information
Year: The specific year(s) or period of data collection. Since many datasets are published months or years after their collection, reporting the actual collection date is crucial for understanding the captured threat landscape and traffic characteristics (e.g., popular applications). For continuously ongoing captures, such as MAWILab [50], we denote the value as Now.
@@
We highlight these three traffic origins due to differences in data characteristics, most notably realism. In general, real-world traffic is considered realistic as it offers genuine network interactions. However, it suffers from mentioned privacy concerns and may lack coverage of specific scenarios, such as flash crowd events [53], if they are not present during the capture.
@@
While traffic generation methods significantly impact realism, the environment type also plays a role. In general, physical deployments for traffic simulation are considered more realistic than virtual ones [55]. Therefore, purely virtual emulation is expected to produce less realistic results than in physical environments, although the differences might be minor.
Naturally, synthetic traffic created without any network activity is considered the least realistic. However, its high flexibility and lack of privacy concerns make it a promising research direction to tackle highly dynamic and ever-changing environments such as intrusion detection. Synthetic traffic has already shown promising results in generating normal traffic with statistically similar properties to more realistic data [56]. Section 5.5.2 discusses trends in traffic generation in more detail.
@@
The attack origin environment is defined the same as for normal traffic – real, emul-{p, v, pv}, and synth. Real attacks correspond to traffic with actual malicious intent captured in real-world production networks, i.e., by detecting and sampling an ongoing attack, as in the Booters dataset [57]. Emulated attacks are launched in controlled environments just to simulate attacks’ behavior, usually using the same tools as attackers (e.g., in Kali Linux [58]). Finally, synthetic attacks were not captured in a network but rather generated by specialized algorithms.
Although in between real and emulated categories, we classify traces of sandboxed malware connected to the Internet (e.g., CTU-13 [59]) as real traffic, assuming its source code was not altered. In this case, we consider its patterns representative of real malware (e.g., spreading, Command & Control server communication) even when executed in a controlled environment.
@@
For instance, the string emul-v (10) DRO for the Bot-IoT dataset [63] indicates 10 different attack classes generated by emulation in a virtualized environment. They belong to (D)DoS, Reconnaissance, and Other attack categories, as the dataset contains traces of data theft and keylogging.
@@ 4.2 Nature of Data
  • packet: Data provided in a per-packet manner. Most such datasets contain raw (binary) packets distributed as Packet Capture (PCAP) files. However, some (e.g., AWID2 [65]) rather provide already pre-extracted full packets or their specific fields in a textual format. This survey distinguishes them by the packet (bin) and packet (txt) identifiers.
@@
  • other: Specific data formats that do not correspond to any of the above. An example of this case is UNR-IDD [77], which provides feature vectors as per-port statistics.
Table 3: Common features for uni-directional flow datasets. Note that bi-directional datasets would include separate fields for features 8 – 11, tracking source $\rightarrow$ destination and destination $\rightarrow$ source traffic as separate entries, resulting in 15 features. Additional features may also be added for data labels. Identifiers and some features might vary across datasets and protocols – e.g., NetFlow v9 [66] uses FIRST_SWITCHED as the first flow timestamp and LAST_SWITCHED for the last instead of defining duration explicitly. Nevertheless, despite naming differences, feature semantics remain identical across datasets.
@@
Anonymized: In some cases, such as captures from real-world networks, the data require anonymization to preserve privacy or hide network details. This property lists parts of the dataset its authors decided to anonymize. While anonymizing IP addresses via prefix-preserving algorithms like CryptoPan [81] is typically sufficient for flow-based data, full packet captures require additional actions. Oftentimes, MAC addresses are also obscured, and packet payloads trimmed to a certain length (e.g., to retain application headers) or removed entirely.
@@ 4.3 Data Volume
A special case of data continuity is periodicity (periodic), which refers to continuous blocks recurring in a regular pattern, such as specific hours per day or days per week (e.g., weekdays). Periodic captures are noted as periods x duration. For instance, the DARPA 1998 dataset [82] was captured over nine weeks on weekdays. Since it also contains a train-test split, it is listed as 7x5d + 2x5d periodic, indicating seven five-day periods for training and two for testing, with each period being continuous.
@@ 4.4 Network Properties
Complete capture: Specifies whether the dataset contains data from the entire underlying network. In general, it is desirable to capture all network communication to ensure traffic diversity and realism. However, some datasets capture only from a single host (e.g., honeypot) or a specific set of hosts due to privacy or practical reasons. Such a setup might skew the traffic patterns (e.g., honeypots would receive significantly more malicious traffic than regular hosts [83]). Therefore, specialized evaluation processes or NIDS types should be employed when dealing with incomplete captures.
@@ 4.5 Evaluation
Split: Defines whether the dataset offers pre-split train and test subsets for evaluation. This is desirable, as it allows transparent comparisons of different NIDSs evaluated on the same data. On the other hand, datasets without a train-test split have to be manually split by their users, potentially leading to evaluation bias and selective reporting [15, 84].
Labels: Specifies whether the dataset is labeled – a crucial property in analyzing and comparing NIDSs. We assign the value yes only if an unambiguous mapping between data instances and classes exists. This applies to direct labels in feature vectors but also to cases when classes can be separated via filters (e.g., IP-based packet labeling in 2017-SUEE8 [85]) or directory structures (e.g., CAIDA DDoS 2007 [86]).

### v3

@@ Abstract
[inst2] organization=Kempelen Institute of Intelligent Technologies, addressline=Sky Park Offices, Bottova 7939/2A, city=Bratislava, postcode=811 09, country=Slovakia
@@ 1 Introduction
The only truly secure system is one that is powered off, cast in a block of concrete and sealed in a lead-lined room with armed guards – and even then I have my doubts.[57]. The legendary quote by Eugene H. Spafford from 1989 is not any less relevant nowadays. Each year, cyberattacks like malware or denial of service (DoS) cause losses of trillions of USD [68]. With recent attacks on hospitals and medical devices, even people’s lives are in danger [163]. Therefore, robust security is required to minimize an attack surface and mitigate potential harm.
@@
An Intrusion Detection System (IDS) is a device or software application monitoring the activity of a target system or a computer network to identify their unauthorized use, misuse, or abuse [150]. Depending on the attack detection paradigm, we distinguish between signature-based (react on known attack patterns) and anomaly-based (react on unknown deviation from the norm) systems. Such systems can operate on the host level – Host-based IDS (HIDS) or the network – Network-based IDS (NIDS). Additional IDSs characteristics include the ability to operate in real-time and whether they merely report the attack or actively participate in its mitigation [215, 33].
@@
Despite the dire need for data, a general consensus within the network intrusion detection (NID) domain over the past decade was that there had been a lack of quality datasets for evaluating NIDS proposals [2, 134, 199]. This factor has significantly contributed to the poor adoption of ML-based methods into real-world systems despite huge scientific efforts [200].
@@
As we show in this work, data scarcity may no longer be the primary issue. Instead, we argue that datasets exist, but the community is largely unaware of their existence and lacks knowledge of best practices for their creation and usage. This can be largely attributed to the lack of standardized NID data repositories and the neglect of domain-specific properties, leading to biased evaluations and over-optimistic results presented in many recent studies [20].
@@
    In order to aid in data selection, we extract 13 properties for each dataset by manually downloading and analyzing the data. We also study the popularity of datasets in state-of-the-art NIDS research. To enhance accessibility, we share our analysis notebooks on GitHub [^1] https://github.com/xGoldy/nid-datasets, allowing others to review dataset structures without downloading the full data.
@@ 2 Survey Methodology
In order to ensure comprehensive, unbiased, and reproducible results, we followed the Systematic Literature Review (SLR) methodology for computer science [105]. This section outlines the survey’s scope and research questions, with further details provided in A.
@@ 3. RQ3:
    Regarding the data properties from RQ2, are there any prevalent patterns or trends in dataset creation?
@@ 3 Domain-Specific Properties and Their Impact on Data
The research on intrusion detection (ID) initially revolved around static signatures or simple statistical methods. Nevertheless, they require significant manual interventions, such as maintaining signature databases and threshold values. Motivated to minimize the need for manual labor and scale to big data more efficiently, the research community has gradually adopted data-driven approaches for ID. Nowadays, AI, and especially ML, are crucial parts of ID research [16].
@@ 3.1 When Hell Comes to Earth: Unpleasant Properties of the Network Intrusion Detection Domain
Network intrusion detection combines properties of both cyberthreat detection and computer networking domains. These properties significantly influence how the data are collected, handled, and interpreted. If machine learning is used for detection, an assumption on the data to be independent and identically distributed (i. i. d.) is also introduced [62].
In essence, the i. i. d. assumes the data to be uncorrelated, while the data for training a model to be similar, i.e., drawn from the same distribution, as the future observed data. If not met, the model’s performance might be negatively affected. However, the intrinsic properties of cyberthreat detection and computer networking interfere with this assumption and thus impair ML deployment in real-world scenarios [16]. In the following paragraphs, we dive deeper into these properties and outline their potential impacts, as summarized in Figure 1.
@@
    Adversarial environment: In addition to natural networks’ dynamicity, the NID domain has to assume the implicit presence of adversaries. Due to the endless cybersecurity arms race, attacks continuously evolve, and new vulnerabilities are discovered. Moreover, the attackers may also attempt to bypass detection via adversarial samples crafted with obfuscation techniques or tiny perturbations of the input data [89].
@@
    High cost of errors: Unlike other ML applications, the relative cost of any output error (i.e., misclassification) is very high. Regular applications can optimize for lower false alarm (false positive) or higher detection (true positive) rates. However, in NID, both need to be considered. Even a single missed anomaly might lead to a successful intrusion with dire consequences. On the other hand, a low relative false alarm rate can still create a tremendous amount of actual alarms due to a traffic imbalance, known as base-rate fallacy [21]. For instance, with a 99% true positive and 1% false positive rate in a 1:100 class balance, we receive 100 false alarms for every 99 true attacks.
@@
    Data confidentiality: Captures from real networks might contain sensitive and confidential information, such as IP addresses or packet payloads, significantly limiting their sharing. Although anonymization techniques [47], like IP address remapping or payload stripping, can be used, they decrease the data’s value – e.g., disabling payload-analysis NIDSs.
@@ 3.2 Limitations of Existing Datasets
    Testbed limitations: Due to the limitations of real-world data, most public datasets contain emulated traffic captured in a lab testbed environment. Despite some benefits of such an approach, its data are not guaranteed to resemble realistic network behavior and threats [147]. Such datasets tend to model over-simplified network conditions, failing to capture their real-world complexity and diversity [38, 73]. The lack of realism in benchmark data is considered a major obstacle in deploying anomaly-based NIDSs in real-world production environments [200, 222]. For instance Catillo et al. [37] have shown that many attacks from DoS public datasets do not resemble real-world characteristics due to being inefficient (not achieving their malicious intent) or assuming non-realistic environment configurations.
@@
    Reproducibility: In computer networking, it is infeasible to collect exactly the same data twice, even in identical environments. This property disallows the reproducibility of the raw packet data. Although it would not be a critical limitation by itself, reproducibility is further restricted for datasets publishing only extracted features (e.g., network flows) without packet data. Even if some datasets do provide raw data, they rarely publish source code and configuration for feature extraction and labeling. In all cases, reproducing the dataset (or even extracting the same features from packets) is infeasible, disallowing data validation and its possible extensions [115].
@@
    Incorrect labeling: Due to networking dynamicity and the common methods for traffic labeling (based on the attack time and the source and destination addresses of the attacker/victim [83]), NID data are prone to mislabeling. The issue has been demonstrated by Liu et al. [127] and Flood et al. [73] through analyses of several popular benchmark datasets. Such mislabeling can result in unrealistic results and benchmark destabilization [158].
@@
    Methodically wrong data handling: In addition to mislabeling, several other factors, primarily originating from human error in data collection and processing, affect the final dataset. A common issue we encountered was improper timestamping, as some datasets contain incorrect timestamps or time inconsistencies of events (e.g., attacks) with the documentation. We also observed disordered packets, undocumented capture gaps, and capture interruptions, resulting in improperly terminated files. Data processing issues, often due to bugs in feature extractor software [38], include missing values, duplicated features or entries, and zero-variance features (e.g., all zeros). These issues are frequently unreported, creating ambiguity for users and undermining the soundness of their conclusions.
@@ 4 Network Intrusion Data and Its Properties
Depending on the dataset goal and capture attributes, we characterize network intrusion data by various properties that influence their use cases and the soundness of methods utilizing them. In this survey, we collected 13 essential properties to provide an unbiased, comprehensive view and comparison of available datasets for NID purposes. A portion of them was adopted from Ring et al. [178]’s survey.
@@ 4.1 General Information
Year: The specific year(s) or period of data collection. Since many datasets are published months or years after collection, reporting the actual collection date is crucial for understanding the captured threat landscape and traffic characteristics (e.g., popular applications). For continuously ongoing captures, such as MAWILab [74], we denote the value as Now.
@@
We highlight these three traffic origins due to differences in data characteristics, especially realism. In general, real-world traffic is considered ideal as it offers genuine network interactions. However, it suffers from mentioned privacy concerns and may lack coverage of specific events, such as flash crowds [19], if they are not present during the capture.
@@
While traffic generation methods significantly impact realism, the environment type also plays a role. In general, physical deployments for traffic simulation are considered more realistic than virtual ones [42]. Therefore, purely virtual emulation is expected to produce less realistic results than physical, although the differences might be minor.
Naturally, synthetic traffic created without any network activity is considered the least realistic. However, its high flexibility and lack of privacy concerns make it a promising research direction to tackle highly dynamic and ever-changing domains such as intrusion detection. Synthetic traffic has already shown promising results in generating normal traffic with statistically similar properties to more realistic data [188]. Section 5.5.2 discusses trends in traffic generation in detail.
@@
The attack origin environment is defined the same as for normal traffic – real, emul-{p, v, pv}, and synth. Real attacks correspond to traffic with actual malicious intent captured in real-world production networks, i.e., by detecting and sampling an ongoing attack, as in the Booters dataset [182]. Emulated attacks are launched in controlled environments just to simulate attacks’ behavior, usually using the same tools as real cyber-attackers (e.g., in Kali Linux [^3] Kali Linux is a Linux distribution designed for digital forensic analysis and penetration testing. Homepage: https://www.kali.org/.). Finally, synthetic attacks were not captured in a network but rather generated by specialized algorithms.
Although in between real and emulated categories, we classify traces of sandboxed malware connected to the Internet (e.g., CTU-13 [78]) as real traffic, assuming its source code was not altered. In this case, we consider its patterns (e.g., spreading, Command & Control traffic) representative of real malware even when executed in a controlled environment.
@@
For instance, the string emul-v (10) DRO for the Bot-IoT dataset [110] indicates 10 different attack classes generated by emulation in a virtualized environment. They belong to (D)DoS, Reconnaissance, and Other attack categories due to data theft and keylogging traces.
@@ 4.2 Nature of Data
  • packet: Data provided in a per-packet manner. Most such datasets contain raw (binary) packets distributed as Packet Capture (PCAP) files. However, some (e.g., AWID2 [107]) provide already pre-extracted full packets or their specific fields in a textual format. We distinguish them by the packet (bin) and packet (txt) identifiers.
@@
  • other: Specific data formats not corresponding to any of the above. An example of this case is UNR-IDD [53], which provides feature vectors as per-port statistics.
Table 3: Common features for uni-directional flow datasets. Note that bi-directional datasets include separate fields for features 8–11, tracking source $\rightarrow$ destination and destination $\rightarrow$ source traffic separately, resulting in 15 features. Additional features may be added for data labels. Identifiers and some features might vary across datasets and protocols – e.g., NetFlow v9 [44] uses FIRST_SWITCHED as the first flow timestamp and LAST_SWITCHED for the last instead of defining duration explicitly. Nevertheless, despite naming differences, feature semantics remain identical across datasets.
@@
Anonymized: In some cases, such as captures from real-world networks, the data require anonymization to preserve privacy or hide network details. This property lists parts of the data its authors decided to anonymize. While anonymizing IP addresses via prefix-preserving algorithms like CryptoPan [227] is typically sufficient for flow-based data, full packet captures require additional actions. Oftentimes, MAC addresses are also obscured, and packet payloads trimmed to a certain length (e.g., to retain application headers) or removed entirely.
@@ 4.3 Data Volume
A special case of continuity is periodicity (periodic), which refers to continuous blocks recurring in a regular pattern, such as specific hours per day or days per week. Periodic captures are noted as periods x duration. For instance, the DARPA 1998 dataset [124] was captured over nine weeks on weekdays. Since it also contains a train-test split, it is listed as 7x5d + 2x5d periodic, indicating seven five-day periods for training and two for testing, with each period being continuous.
@@ 4.4 Network Properties
Complete capture: Specifies whether the dataset contains data from the entire underlying network. In general, it is desirable to capture all network communication to ensure traffic diversity and realism. However, some datasets capture only from a single host (e.g., honeypot) or a specific set of hosts due to privacy or practical reasons. Such a setup might skew the traffic patterns (e.g., honeypots would receive significantly more malicious traffic than regular hosts [190]). Therefore, specialized evaluations or NIDS types should be employed when dealing with incomplete captures.
@@ 4.5 Evaluation
Split: Defines whether the dataset offers pre-split train and test subsets. This is desirable, as it allows transparent comparisons of different NIDSs evaluated on the same data. On the other hand, datasets without a train-test split have to be split by their users, potentially leading to evaluation bias and selective reporting [20, 126].
Labels: Specifies whether the dataset is labeled – a crucial property in analyzing and comparing NIDSs. We assign the value yes only if an unambiguous mapping between data instances and classes exists. This applies to direct labels in feature vectors but also when classes can be separated via filters (e.g., IP-based packet labeling in 2017-SUEE8 [130]) or directory structures (e.g., CAIDA DDoS 2007 [34]).